# Graph-theoretical estimates of the diameters of the Rubik’s Cube groups

### v1

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order ($n$) and other local parameters of the graph such as the degree, even girth $g$ ($\geq 4$), and number of cycles of length $g$ passing through a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of the Rubik’s Cube groups of various sizes and metrics, yielding reasonably tight lower bounds, which range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s number.Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014a, 2014b) It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. Previously,Hirata (2024) we estimated these diameters for the Cubes of various sizes in different metrics (the set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph. For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notationSingmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ (Fig. 1). The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metricSingmaster (1981) ($R^{2}$, $D^{2}$, and $B^{2}$) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2).
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. The order ($n$) is 8, the girth ($g$) is 4, the diameter ($d$) is 3, and the number of cycles with length $g$ per vertex ($\eta$) is 3.
@@
Figure 2: Generalized Petersen graph $G(12,5)$ as the Cayley graph of the $2\times 2\times 2$ Rubik’s Cube group in the square-turn metric. The order ($n$) is 24, the girth ($g$) is 6, the diameter ($d$) is 4, and the number of cycles with length $g$ per vertex ($\eta$) is 3.
@@
The Cayley graphs of the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cube groups in the quarter-turn metrics should also be symmetric graphs because all configurations and all turns are symmetrically equivalent. The Cayley graph of the 3$\times$3$\times$3 Cube group in the square metric is another symmetric graph, which turns out to be locally isomorphic to the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric, with the only difference being their orders (numbers of vertexes). In this article, we introduce the formula for a strict lower bound for the diameter of a general symmetric graph derived with algebraic-graph-theoretical arguments.Biggs (1993); West (2019) It can be evaluated with the order and some easily determined local parameters of the graph, such as the degree, girth ($g$), and number of cycles with length $g$ passing through a vertex. We then apply this formula to the diameters of the Cube groups of different sizes and metrics and compare the lower bounds with the probabilistic estimatesHirata (2024) and actual values.
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k$, let $g$ ($\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ passing through a given vertex. We seek a lower bound for $d$, whose precise determination is generally hard, given in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available.
@@
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any metric that does not form a mathematically well-defined group. It has been derived on the basis of a probabilistic argument or a modified version of the coupon collector’s problem. Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata, 2024, a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (LABEL:dmin1)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21).
@@ IV Cayley graphs of Rubik’s Cube groups
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of cycles with length $g$ per vertex. $r_{\text{max}}$ [^2] The maximum branching ratio at distance $g/2-1$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al.Rokicki et al. (2014b, 2013, a) Other diameters listed can be determined more easily. $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (LABEL:dmin1) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)].Hirata (2024)
3$\times$3$\times$3 Square-slice $8$ $3$ $4$ $3$ $1.0$ $3$ $3$infty
2$\times$2$\times$2 Square $24$ $3$ $6$ $3$ $1.5$ $4$ $4$ $10.0$
2$\times$2$\times$2 Quarter $3.67\times 10^{6}$ $6$ $4$ $3$ $4.5$ $14$ $10$ $13.4$
3$\times$3$\times$3 Sqaure $6.63\times 10^{5}$ $6$ $4$ $3$ $4.5$ $15$ $9$ $11.9$
3$\times$3$\times$3 Quarter $4.33\times 10^{19}$ $12$ $4$ $18$ $9.5$ $26$ $20$ $24.8$
@@ IV.1 The 3$\times$3$\times$3 Cube in the square-slice-turn metric
The order of the graph $n$ is a global parameter and is potentially hard to determine. However, for the Cayley graph of a Rubik’s Cube group, it can be worked out by a careful combinatorial argumentHofstadter (1981); Singmaster (1981); Hirata (2024) or with the aid of group theory.Joyner (2008) In the square-slice metric,Singmaster (1981) each of the three center slices is allowed to independently take one of two orientations and thus $n=2^{3}=8$. Other local parameters such as the degree $k$, girth $g$, and the number $\eta$ of cycles of length $g$ per vertex can be gleaned from the small neighborhood of a vertex in the partial graph. Inspecting Fig. 1, we find that $k=3$, $g=4$, and $\eta=3$.
The distance array is $\bm{d}=\{1,3,3,1\}$, which obeys Eqs. (6) and (8), as it must. The branching ratios are $\bm{r}=\{3,1,1/3\}$, satisfying Eq. (11).
Using Eq. (14), $r_{\text{max}}=1$, which corresponds to the predicted distance array of $\bm{d}=\{1,3,3,3\}$. Equation (20) yields $d_{\text{min}}=\lceil 2.3\rceil=3$, which is a tight lower bound for the actual diameter of 3. The probabilistic estimate [Eq. (21)] breaks down in this tiny case because $n=8$ is too few for probabilistic logic.
@@ IV.2 The 2$\times$2$\times$2 Cube in the square-turn metric
Substituting these values into Eq. (LABEL:dmin1), we obtain $d_{\text{min}}=\lceil 3.4\rceil=4$, which is a tight lower bound for the actual diameter of 4 (Table 1).
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Figure 3: A partial Cayley graph of the 2$\times$2$\times$2 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree ($k$) 6. The order ($n$) is $3.67\times 10^{6}$, the girth ($g$) is 4, the diameter ($d$) is 14, and the number of cycles with length $g$ per vertex ($\eta$) is 3. The vertexes on the periphery are further connected to other vertexes that are not shown.
@@
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, the Cayley graph of this group should also be edge-transitive and thus symmetric.Biggs (1993); West (2019) This is supported by the appearance of the local graph of Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. A smallest portion of the Cayley graph is sufficient to determine that $k=6$, $g=4$, and $\eta=3$.
@@
Using Eq. (LABEL:dmin1), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected.
@@ IV.4 The 3$\times$3$\times$3 Cube in the square-turn metric
Figure 5: A partial Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the square-turn metric, which is symmetric with degree ($k$) 6. The order ($n$) is 663 552663552663\,552663 552, the girth ($g$) is 4, the diameter ($d$) is 15, and the number of cycles with length $g$ per vertex ($\eta$) is 3.
The square-turn metric of the 3$\times$3$\times$3 Rubik’s Cube allows six 180 rotations denoted by $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, and $F^{2}$ in the Singmaster notation.Singmaster (1981) The order of the Cayley graph can be determined by group theory to be 663 552663552663\,552663 552.Joyner (2008) Combinatorial logic explains the same number as follows:
@@
The first two factors $4!4$ is the number of ways in which the eight corner cubies can be positioned with their orientations fixed by the positions. They can be viewed as the four diagonal corner cubies of the 2$\times$2$\times$2 Cube in the square metric, which has $4!=24$ different configurations, which is further multiplied by the 4 possible positions of the ulf cubie (which was held fixed in the 2$\times$2$\times$2 Cube in the square metric, but now can move) reachable by the $L^{2}$, $U^{2}$, and $F^{2}$ turns. The remaining 3 corner cubies’ positions and orientations are completely determined by the first five. The quotient $4!4!4!/2$ is the number of ways three groups of 4 edge cubies (sharing a common center slice) can be positioned with their orientations fixed. Since the number of correctly oriented edge cubies is always even, there are two orbits and the number must be divided by two. In short, the global parameter $n$ can be known analytically.
What makes this group interesting is the fact that its Cayley graph is locally isomorphic with the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric (Sec. IV.3). The only difference is the orders of these graphs; one is 5.5 times smaller than the other. Figure 5 shows the local (up to distance 3) Cayley graph, which is indeed identical with Fig. 3; it is a symmetric graph with $k=6$, $g=4$, and $\eta=3$. Surprisingly, the diameter ($d=15$) of this graph with $n=6.63\times 10^{5}$ is greater than the diameter ($d=14$) of the 2$\times$2$\times$2 Cube group in the quarter-turn metric with $n=3.67\times 10^{6}$. With all the other parameters being the same, it is impossible to correctly predict this inversion by the simple graph-theoretical or probabilistic estimations presented here and earlier.Hirata (2024) These two graphs, therefore, underscore the diversity of symmetric graph structures differing only in their orders.
Using Eq. (LABEL:dmin1), we obtain as the lower bound of this graph $d_{\text{min}}=9$, which is lower than $d_{\text{min}}=10$ of the 2$\times$2$\times$2 Cube group in the quarter-turn metric. Likewise, the probabilistic estimate $d_{\text{probab}}=11.9$ is also lower than that ($d_{\text{probab}}=13.4$) of the 2$\times$2$\times$2 Cube group in the quarter-turn metric and is uncharacteristically underestimated. Hence, the inversion is not reproduced.
Figure 4 compares the distance arrays $\bm{d}$ of these two locally isomorphic Cayley graphs. The two curves initially rise at the same rate for they have the same $r_{\text{max}}$. The curve for the 3$\times$3$\times$3 Cube group in the square metric (the red open circles) begins to slow the growth earlier, which is expected because its $n$ is 5.5-times smaller. The peak in $\bm{d}$ occurs slightly earlier (by one half distance unit) in the 3$\times$3$\times$3 Cube group in the square metric with the smaller $n$. What causes the inversion is the decay phase; the $\bm{d}$ of the 3$\times$3$\times$3 Cube group in the square metric decays unusually slowly (see other plots of $\bm{d}$ in Ref. Hirata, 2024 also), considerably delaying reaching the diameter. The cause of this slow decay is unknown.

### v2

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order ($n$) and other local parameters of the graph such as the degree, even girth $g$ ($\geq 4$), and number of $g$-cycles traversing a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of the Rubik’s Cube groups of various sizes and metrics, yielding slightly tighter lower bounds than those of random $k$-regular graphs proposed by Bollobás and de la Vega, which range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. Each vertex is a unique configuration (a pattern of face colors of the cubies while holding the orientation of the whole Cube fixed) reachable by a combination of the $R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ turns (in the Singmaster notationSingmaster (1981)) from the solved configuration. Each edge connecting two vertexes is a turn (and its inverse) that brings one configuration denoted by one of the vertexes to another represented by the other vertex. The order ($n$) of the graph is 8, the girth ($g$) is 4, the diameter ($d$) is 3, and the number ($\eta$) of $g$-cycles traversing a vertex is 3.
@@
Figure 2: Generalized Petersen graph $G(12,5)$ as the Cayley graph of the $2\times 2\times 2$ Rubik’s Cube group in the square-turn metric. The order ($n$) is 24, the girth ($g$) is 6, the diameter ($d$) is 4, and the number ($\eta$) of $g$-cycles per vertex is 3.
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s number.Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014a, 2014b) It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. Previously,Hirata (2024) we estimated these diameters for the Cubes of various sizes in different metrics (a set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph.Biggs (1993); West (2019) For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notationSingmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ shown in Fig. 1.West (2019) The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metricSingmaster (1981) ($R^{2}$, $D^{2}$, and $B^{2}$) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2).West (2019)
@@
The Cayley graphs of the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cube groups in the quarter-turn metrics should also be symmetric graphs because all configurations and all turns are symmetrically equivalent. The Cayley graph of the 3$\times$3$\times$3 Cube group in the square metric is another symmetric graph, which turns out to be locally isomorphic to the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric, with the only difference being their orders (numbers of vertexes).
In this article, we introduce the formula for a strict lower bound for the diameter of a general symmetric graph derived with algebraic-graph-theoretical arguments.Biggs (1993); West (2019) It can be evaluated with the order and some easily determined local parameters of the graph, such as the degree, girth ($g$), and number of $g$-cycles traversing a vertex. We then apply this formula to the diameters of the Cube groups of different sizes and metrics and compare the lower bounds with the probabilistic estimatesHirata (2024) and actual values. See Ref. Hirata, 2024 for nomenclature of the Rubik’s Cube groups.
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k$, let $g$ ($\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ traversing a vertex. We seek a lower bound for $d$, whose precise determination is generally hard, given in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available.
@@
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any metric that does not form a mathematically well-defined group. It has been derived on the basis of a probabilistic argument or a modified version of the coupon collector’s problem.Hirata (2024) Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata, 2024, a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (LABEL:dmin1)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21); they become nearly identical in the $n\to\infty$ limit.
Upper and lower bounds for the diameter of random $k$-regular graphs were obtained by Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001) For “almost every” graph and a constant $\epsilon>0$, the BV bounds are
$d_{\text{min}}^{\text{BV}}$ $=$ $\left\lfloor\log_{k-1}n\right\rfloor+\left\lfloor\log_{k-1}\ln n-\log_{k-1}\frac{6k}{k-2}\right\rfloor+1,$ (22)
$d_{\text{max}}^{\text{BV}}$ $=$ $\left\lceil{\log_{k-1}\left\{(2+\epsilon)\,kn\ln n\right\}}\right\rceil+1,$ (23)
asymptotically ($n\to\infty$). These are derived on the basis of probabilistic arguments not unlike the one leading to Eq. (21). Hence, the lower bound [Eq. (22)] bears resemblance with the first term of the probabilistic estimate [Eq. (21)] or $d_{\text{min}}$ [Eq. (LABEL:dmin1)]. As $n\to\infty$, the first term of the former dominates and approaches the latter from below because $r_{\text{max}}\leq k-1$ [see Eq. (14)]. In what follows, the BV bounds are computed with $\epsilon=0$.
@@ IV Cayley graphs of Rubik’s Cube groups
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of $g$-cycles per vertex. $r_{\text{max}}$ [^2] Upper bound for the branching ratio at distance $g/2-1$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al.Rokicki et al. (2014b, 2013, a) Other diameters listed are also determined computationally more easily. $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (LABEL:dmin1) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)].Hirata (2024) $d_{\text{min}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001) $d_{\text{max}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001)
3$\times$3$\times$3 Square-slice $8$ $3$ $4$ $3$ $1.0$ $3$ $3$infty 00 $8$
2$\times$2$\times$2 Square $24$ $3$ $6$ $3$ $1.5$ $4$ $4$ $10.0$ $2$ $10$
2$\times$2$\times$2 Quarter $3.67\times 10^{6}$ $6$ $4$ $3$ $4.5$ $14$ $10$ $13.4$ $10$ $14$
3$\times$3$\times$3 Sqaure $6.63\times 10^{5}$ $6$ $4$ $3$ $4.5$ $15$ $9$ $11.9$ $9$ $13$
3$\times$3$\times$3 Quarter $4.33\times 10^{19}$ $12$ $4$ $18$ $9.5$ $26$ $20$ $24.8$ $19$ $23$
@@ IV.1 The 3$\times$3$\times$3 Cube in the square-slice-turn metric
The order of the graph $n$ is a global parameter and is potentially hard to determine. However, for a Rubik’s Cube group, it can be worked out by a careful combinatorial argumentHofstadter (1981); Singmaster (1981); Hirata (2024) or with the aid of group theory.Joyner (2008) In the square-slice metric,Singmaster (1981) each of the three center slices is allowed to independently take one of two orientations and thus $n=2^{3}=8$. Other local parameters such as the degree $k$, girth $g$, and the number $\eta$ of $g$-cycles per vertex can be gleaned by inspecting a small neighborhood of a vertex in the partial graph. From Fig. 1, we find that $k=3$, $g=4$, and $\eta=3$.
The distance array is $\bm{d}=\{1,3,3,1\}$, which obeys Eqs. (6) and (8) as it must. The branching ratios are $\bm{r}=\{3,1,1/3\}$, satisfying Eq. (11).
Using Eq. (14), $r_{\text{max}}=1$, which corresponds to the predicted distance array of $\bm{d}=\{1,3,3,3\}$. Equation (20) yields $d_{\text{min}}=\lceil 2.3\rceil=3$, which is a tight lower bound for the actual diameter of 3. The probabilistic estimate [Eq. (21)] breaks down in this tiny case because $n=8$ is too few for probabilistic logic. For the same reason, the BV lower and upper bounds [Eqs. (22) and (23)] are too low and too high to be useful.
@@ IV.2 The 2$\times$2$\times$2 Cube in the square-turn metric
Substituting these values into Eq. (LABEL:dmin1), we obtain $d_{\text{min}}=\lceil 3.4\rceil=4$, which is a tight lower bound for the actual diameter of 4 (Table 1). This graph is still too small for probabilistic arguments and $d_{\text{probab}}=d_{\text{max}}^{\text{BV}}=10$ are too high, while $d_{\text{min}}^{\text{BV}}=2$ is too low.
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Figure 3: A partial Cayley graph of the 2$\times$2$\times$2 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree ($k$) 6. The order ($n$) is $3.67\times 10^{6}$, the girth ($g$) is 4, the diameter ($d$) is 14, and the number ($\eta$) of $g$-cycles per vertex is 3. The vertexes on the periphery are further connected to other vertexes that are not shown.
@@
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, the Cayley graph of this group should also be edge-transitive and thus symmetric.Biggs (1993); West (2019) This is supported by the appearance of a partial graph (within distance 3 from an arbitrarily chosen origin vertex) drawn in Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. This smallest portion of the Cayley graph is sufficient to determine $k=6$, $g=4$, and $\eta=3$.
@@
Using Eq. (LABEL:dmin1), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected. The BV lower bound $d_{\text{min}}^{\text{BV}}=10$ [Eq. (22)] is the same as $d_{\text{min}}$, which is not surprising considering the similarity in their underlying logic. The $d_{\text{max}}^{\text{BV}}=14$ is a tight upper bound and equal to the actual diameter ($d=14$) itself, but this is likely the result of error cancellation (see below). On the other hand, the good agreement between $d_{\text{probab}}=13.4$ and the actual $d$ is not accidental.Hirata (2024)
@@ IV.4 The 3$\times$3$\times$3 Cube in the square-turn metric
Figure 5: A partial Cayley graph of the 3$\times$3$\times$3 Rubik’s Cube group in the square-turn metric, which is symmetric with degree ($k$) 6. The order ($n$) is 663 552663552663\,552663 552, the girth ($g$) is 4, the diameter ($d$) is 15, and the number ($\eta$) of $g$-cycles per vertex is 3.
The square-turn metric of the 3$\times$3$\times$3 Rubik’s Cube allows six 180 rotations denoted by $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, and $F^{2}$ in the Singmaster notation.Singmaster (1981) The order of the Cayley graph can be determined by group theory to be 663 552663552663\,552663 552.Joyner (2008); Hirata (2024)
What makes this group interesting is the fact that its Cayley graph is locally isomorphic with the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric (Sec. IV.3). By “locally isomorphic,” we mean that their partial graphs within distance $g/2$ from an arbitrary origin vertex are pairwise isomorphic. The only difference is the orders of these graphs; one is 5.5 times smaller than the other. Figure 5 shows the local (within distance 3) Cayley graph, which is indeed identical with Fig. 3; it is a symmetric graph with $k=6$, $g=4$, and $\eta=3$. Surprisingly, the diameter ($d=15$) of this graph with $n=6.63\times 10^{5}$ is greater than the diameter ($d=14$) of the 2$\times$2$\times$2 Cube group in the quarter-turn metric with $n=3.67\times 10^{6}$. With all the other parameters being the same, it is impossible to correctly predict this inversion by the simple graph-theoretical or probabilistic estimations presented here and earlier.Hirata (2024); Bollobás and de la Vega (1982); Bollobás (2001) These two graphs, therefore, underscore the diversity of symmetric graph structures differing only in their orders.
@@
Figure 4 compares the distance arrays $\bm{d}$ of these two locally isomorphic Cayley graphs. The two curves initially rise at the same rate for they have the same $r_{\text{max}}$. The curve for the 3$\times$3$\times$3 Cube group in the square metric (the red open circles) begins to slow the growth earlier, which is expected because its $n$ is 5.5-times smaller. The peak in $\bm{d}$ occurs slightly earlier (by about one half distance unit) in the 3$\times$3$\times$3 Cube group in the square metric with the smaller $n$. What causes the inversion is the decay phase; the $\bm{d}$ of the 3$\times$3$\times$3 Cube group in the square metric decays unusually slowly (see other plots of $\bm{d}$ in Ref. Hirata, 2024 also), considerably delaying reaching the diameter. The cause of this slow decay is unknown.